# ThinkRec: Thinking-based recommendation via LLM

### v3

@@ Abstract
Recent advances in large language models (LLMs) have enabled more semantic-aware recommendations through natural language generation. Existing LLM for recommendation (LLM4Rec) methods mostly operate in a System 1-like manner, relying on superficial features to match similar items based on click history, rather than reasoning through deeper behavioral logic. This often leads to superficial and erroneous recommendations. Motivated by this, we propose ThinkRec, a thinking-based framework that shifts LLM4Rec from System 1 to System 2 (rational system). Technically, ThinkRec introduces a thinking activation mechanism that augments item metadata with keyword summarization and injects synthetic reasoning traces, guiding the model to form interpretable reasoning chains that consist of analyzing interaction histories, identifying user preferences, and making decisions based on target items. On top of this, we propose an instance-wise expert fusion mechanism to reduce the reasoning difficulty. By dynamically assigning weights to expert models based on users’ latent features, ThinkRec adapts its reasoning path to individual users, thereby enhancing precision and personalization. Extensive experiments on real-world datasets demonstrate that ThinkRec significantly improves the accuracy and interpretability of recommendations .
1 Introduction
@@
Recommendation systems are indispensable in modern digital platforms, enabling users to navigate vast content efficiently [1, 2, 3]. Traditional sequential recommendation methods rely on implicit modeling of user interaction histories, cannot model context and incorporate broader knowledge, which restricts their reasoning and generalization ability. Recent advances in LLMs offer strong contextual comprehension and extensive world knowledge to improve recommendation systems.
Prior solution paradigms of LLM4Rec can be seperated into three categories: (1) item scoring [4, 5], where LLMs answer binary preference questions given user and item context; (2) item generation [6, 7], which maps natural language prompts to item IDs through aligned embedding spaces; and (3) hybrid modeling [8, 9], where a single LLM performs multiple tasks such as pointwise prediction, pairwise ranking, or listwise generation. Although these LLM4Rec methods differ in output formats and representation learning strategies, they fundamentally resemble System 1—the intuitive system—in cognitive science [10]. They tend to match similar items based on click history while overlooking the deeper behavioral logic. This limitation becomes evident in cases such as the one shown in Figure 1(b), where the user’s behavior over time is: dislikes “Dune”, likes “The Three-Body Problem”, and likes “Foundation” (all three are science fiction). Methods that rely on System 1-like intuition tend to infer that the user would also like “Hyperion” (science fiction) simply because it belongs to the same genre. In reality, the user might dislike philosophical or metaphysical themes in fiction, which are prominent in *Hyperion*, making it an unsuitable recommendation. Clearly, if we can leverage the vast knowledge encoded in LLMs and fully activate their reasoning capabilities for user behavior modeling, recommendation performance can be significantly improved.
@@
Figure 1: (a) shows the composition of user interaction data. (b) illustrates previous LLM-based recommendations. (c) illustrates our ThinkRec consists of extraction, grouping, and thinking. (d) compares ThinkRec with baselines on AUC and UAUC.
@@
This motivates our effort to push LLM4Rec from a System 1 paradigm toward a more rational, System 2-like reasoning framework. We raise two key questions for the advancement: 1) How to effectively balance recommendation objectives with language modeling tasks to fully exploit the reasoning capabilities of LLMs. Existing methods often prioritize direct recommendation metrics such as hit rate or ranking accuracy, overlooking LLMs’ strengths in semantic reasoning. However, blind reinforcement thinking can lead to simple next token prediction, defeating the goal of recommendation. 2) How to think more effectively in the presence of diverse user behaviors and underlying preferences. As shown in Figure 1(c), user behaviors vary widely, and uniform modeling tends to obscure personality preferences while introducing noise that distracts the model from salient signals, impairing its reasoning capacity. Moreover, inferring user intent solely from high-rating items and generic world knowledge limits the informational basis for accurate preference reasoning.
@@
To address these challenges, we propose Thinking-based Recommendation via LLM, abbreviated as ThinkRec. One of the main problems faced is that the data and optimization goals of recommendation tasks lack the ability to activate thinking in LLMs (Challenge 1). To overcome this challenge, we designed the item augmentation and thinking activation framework for finetuning. The fine-tuned model analyzes associations in historical item information, determines user preferences, and gives explicit reasons while deriving recommendations. We extracted the metadata keywords of the items with the help of an existing summarization model as the augmentation information of the items to support the reasoning. In addition, reasoning data is synthesized using a strong reasoning model, and the reasoning capability is distilled to the local model by mixed sampling of reasoning and recommendation data. Therefore, item augmentation and thinking activation become a bridge connecting the recommendation task and the language task, making recommendations traceable. To address the difficulty of thinking diversely with rich information (Challenge 2), we add the user’s preferences (yes/no) of items to the prompts and generate personalized recommendation experts based on latent user features. In the technique, we design a dynamic Low-Rank Adaptation (LoRA) fusion method. Users are grouped by latent user features extracted from traditional recommendation models as shown in Figure 1(c). A set of base LoRAs can be represented and fine-tuned using the grouped data, and the engagement level of each LoRA can be determined through a gating mechanism. The difficulty of thinking is reduced through information classification and personalization mechanisms.
@@
We conduct experiments on three real-world recommendation datasets, validating the rationality and effectiveness of ThinkRec. ThinkRec average outperforms state-of-the-art baselines by 7.96% in AUC and by 56.54% in METEOR. In summary, the main contributions of this work are threefold:
  • We analyze the key challenges of LLM-based recommendations and almost the first emphasize the importance of activating thinking to make recommendations more interpretable.
  • We propose the ThinkRec, including item augmentation and a thinking activation framework, as well as a latent user feature-based gating mechanism for instance-wise recommendation expert fusion.
@@
2 Related Work
2.1 LLM-based Recommendation
With the rapid progress of LLMs, there has been growing interest in adapting them to recommendation tasks using both textual and structured data. Early works [11, 12, 13, 5] rely on prompt-based scoring by converting recommendations into binary question-answering, while others [14, 6, 15] align language space with item embeddings to directly generate item IDs. More advanced approaches, such as P5 [16] and InstructRec [9] unify multiple sub-tasks—like rating prediction, pairwise comparison, or ranking, into a single language modeling framework. These methods demonstrate LLMs’ flexibility in expressing recommendation semantics, yet most focus on adapting output formats, treating LLMs as static scorers or selectors with limited interpretability.
@@ 2.1 LLM-based Recommendation
Beyond static prompting, recent efforts shift toward viewing LLMs as interactive agents capable of multi-step reasoning and personalized decision-making. For example, RecMind [17] and MACRec [18] introduce modularized reasoning structures, enabling planning-based recommendation. Others explore integrating LLMs into various stages of the pipeline—for instance, generating user or item representations [19] or acting as high-level decision controllers [20]. Meanwhile, instruction tuning [9], chain-of-thought prompting [21] have been proposed to enhance LLMs’ semantic understanding. Despite these developments, explicit reasoning supervision and fine-grained user modeling remain underexplored. ThinkRec addresses these gaps by integrating reasoning-augmented training and dynamic expert fusion, enabling interpretable, user-specific recommendations grounded in reasoning.
2.2 Reasoning Model
@@
Recent advances have transformed LLMs from passive token predictors into structured reasoning agents by incorporating process-level supervision. Techniques such as Chain-of-Thought prompting [22], ReAct [23], and Tree-of-Thoughts [24] enable models to generate intermediate reasoning steps, improving interpretability and multistep inference quality. To further reinforce reasoning capabilities, recent works have introduced process reward models [25] and self-improvement pipelines using techniques like Monte Carlo Tree Search [26] and reasoning distillation [27]. These trends define the emerging paradigm of Large Reasoning Models, which prioritize explicit, verifiable thinking processes over direct output. ThinkRec draws inspiration from this direction by introducing reasoning-augmented training for recommendation, aligning to embed structured reasoning into decision making.
3 Method
@@
Figure 2: Overview of the proposed ThinkRec. Keywords are extracted from the description of items with the pretrained PolyLM-Qwen-7B. Reasoning data is synthesized with a reasoning model from a sample of training data. Global LoRA is trained with total data, and base LoRAs are tuned on it with data grouping by user feature. The used LoRA, fusioned or single, is determined by the similarity between the user feature and LoRA representations.
3.1 Preliminary
@@
Integrating Collaborative Embeddings into LLMs. The sequential recommenders, such as MF [28] and LightGCN [29], are used for encoding collaborative information, which can be integrated into LLM embeddings. Formally, for each sample $(u,t,i_{id},l)$:
$\mathbf{e}_{s}^{u}=f_{\psi}(u;\mathcal{S});\mathbf{e}_{s}^{i}=f_{\psi}(i_{id};\mathcal{S}),$ (1)
@@
$\mathbf{E}_{txt}=\mathrm{WE}(\mathrm{TKZ}(i_{txt})),$ (2)
where $\mathrm{TKZ}(\cdot),\mathrm{WE}(\cdot)$ present the LLM tokenizer and word embedding layer, respectively. $\mathbf{E}_{txt}=\{\mathbf{emb}_{txt}^{1:L}\}$, where $L$ is the number of tokens and $\mathbf{emb}_{txt}^{1}\in\mathbb{R}^{1\times d_{2}}$ denotes embeddings in the language space with dimension $d_{2}$. To align $\mathbf{e}_{s}^{i},\mathbf{e}_{s}^{u}$ to LLMs’ language space, $proj_{\phi}(\cdot)$ is used:
$\mathbf{emb}_{s}^{i}=proj_{\phi}(\mathbf{e}_{s}^{i}),\mathbf{emb}_{s}^{u}=proj_{\phi}(\mathbf{e}_{s}^{u}).$ (3)
3.2 Thinking Enhanced Recommendation
@@
3.2.1 Reasoning data construction
To support LLMs for more effective reasoning, we need more textual information in addition to item titles. Given that the raw metadata often contains excessive or noisy content, directly using long-form text may hinder the understanding and inference of LLMs. To address this, we leverage a pre-trained summarization model [30] to extract key concepts from the item metadata, and retain up to 10 representative keywords to enhance the semantic representation of each item. To model more comprehensive preference information, not just liking information, we add low-scoring interactions to the history sequence as well. Then we augment each item with title $i_{t}$, label $i_{l}\in\{yes,no\}$, feature, and keywords $i_{k}$ as follows:
$i_{txt}$ = $i_{t}$ with feature $\mathbf{emb}_{s}^{i}$ (evaluation: $i_{l}$) with description: $i_{k}$.
Since existing recommendation datasets lack explicit reasoning traces, we synthesize a subset of samples with explainable rationale by leveraging a powerful reasoning model QwQ [^1] https://qwenlm.github.io/blog/qwq-32b/. Specifically, we sample a few thousand instances from the training data and construct textual prompts according to Appendix B.3. We then repeatedly query the reasoning model until a correct prediction is produced and record the latest explanation content as the reason for interaction as in Figure 2(b).
3.2.2 Thinking activation
@@
$\begin{array}[]{rl}\mathbf{E}^{qa}=&\mathrm{Concat}(\mathbf{E}^{q},\mathbf{E}^{a}),\\ pos=&-\mathrm{Length}(\mathbf{E}^{a}),\end{array}$ (4)
where $\mathbf{E}^{q},\mathbf{E}^{a}$ means embedding of question and answer constructing from $\mathcal{S}$, respectively. While $pos$ marked the position of the first answer token. This mixed sampling setup encourages the model to learn both prediction accuracy and semantically aligned reasoning behavior, within a shared representation space as shown in Figure 2(c).
@@
$\mathcal{L}_{think},logits=\mathrm{LLM}_{\theta}(\mathbf{E}^{qa}).$ (5)
@@
$\begin{array}[]{rl}posid=&\mathrm{TKZ}(\mathrm{{}^{\prime}Yes^{\prime}}),\\ \hat{l}=&logits[pos][posid],\\ \mathcal{L}_{rec}=&\mathrm{BCE}(\hat{l},l).\end{array}$ (6)
To jointly optimize both objectives on a unified batch, we always use two losses, adjusting the weights of each loss according to instances:
=$\mathcal{L}=\left\{\begin{array}[]{rl}\alpha\mathcal{L}_{rec}+\beta\mathcal{L}_{think},&thinking\\ \eta\mathcal{L}_{rec}+\gamma\mathcal{L}_{think},&recommend\end{array}\right..$ (7)
@@
3.3 Recommendation Experts Fusion
3.3.1 Base Expert fine-tuning
To enable each expert to better capture user-specific preferences and behavioral patterns, we partition users into distinct groups based on latent features. Specifically, we utilize user embeddings derived from various pretrained small collaborative models, each of which implicitly encodes the interaction semantics of users. These embeddings serve as the basis for grouping. We aggregate the representations across all users and perform unsupervised clustering to obtain $N$ groups $\mathcal{S}^{\prime}=\{\mathcal{S}_{1:N}\}$. The resulting clusters are then used to partition the training and validating data, allowing each expert to specialize in a subset of users with similar representations to simplify the preference modeling.
We first fine-tune a global expert using the full data under the framework described in Section 3.2.2, applying LoRA-based adaptation to enable generalizable recommendations with thinking. Building on this global expert $\mathrm{LoRA}_{global}$, we further adapt selected LoRA layers (the last 8 layers) using grouped data, enabling the model to preserve general thinking ability while adapting to more nuanced user preferences. As a result, we obtain a set of base experts $\{\mathrm{LoRA}_{1:N}\}$, which serve as candidates for dynamic expert fusion.
3.3.2 Instance Wise Expert Fusion
To determine which expert is most suitable for a given user, we assign the mean of user features extracted within each group by the corresponding small model as representations of experts =$\mathcal{E}=\{\mathbf{e}_{1:N}^{c}\}$. We then compute the match between user features and expert representations to estimate each base expert’s involvement in modeling the user’s preferences. The cosine similarity and softmax functions were used to obtain participation scores $\mathbf{w}^{u}=\{w^{u}_{1:N}\}$ of user $u$ based on experts:
$\begin{array}[]{rl}\mathbf{e}_{n}^{c}=&\mathrm{Mean}(\mathbf{e}^{u}_{s}),u\in\mathcal{S}_{n},\\ \mathbf{w}^{u}=&\mathrm{Softmax}(\mathrm{Cosim(\mathbf{e}_{s}^{u},\mathcal{E})}/\tau),\end{array}$ (8)

### v4

@@ Abstract.
Recent advances in large language models (LLMs) have enabled more semantic-aware recommendations through natural language generation. Existing LLM for recommendation (LLM4Rec) methods mostly operate in a System 1-like manner, relying on superficial features to match similar items based on click history, rather than reasoning through deeper behavioral logic. This often leads to superficial and erroneous recommendations. Inspired by this, we propose ThinkRec, a thinking-based framework that shifts LLM4Rec from an intuitive system to a rational system. First, ThinkRec introduces a thinking activation mechanism by injecting synthetic reasoning traces, making the recommendation process resemble the Chain of Thought (CoT) reasoning of LLMs. This mechanism analyzes interaction histories, identifies user preferences, and makes decisions based on target items. Furthermore, considering the highly diverse distribution of recommendation data, we propose an instance-wise expert fusion mechanism to reduce the reasoning difficulty. By dynamically assigning weights to expert models based on users’ latent features, ThinkRec adapts its reasoning path to individual users, thereby enhancing precision and personalization. Extensive experiments on various real-world web user behavior preference datasets demonstrate that ThinkRec significantly outperforms baselines in terms of recommendation accuracy and interpretability, providing superior recommendations based on a deeper understanding of user intent and a more rigorous reasoning process. Code is available in https://github.com/Yu-Qi-hang/ThinkRec.
LLM, Recommendation System, Reasoning Model
journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesdoi: 10.1145/3774904.3792070isbn: 979-8-4007-2307-0/2026/04ccs: Information systemsccs: Information systems Information retrievalccs: Information systems Retrieval tasks and goalsccs: Information systems Recommender systems
1. Introduction
@@
Recommendation systems are indispensable in modern digital platforms, enabling users to navigate vast content efficiently (Glenski et al., 2021; Ma et al., 2023; Iana et al., 2024). Traditional sequential recommendation methods rely on implicit modeling of user interaction histories and are limited in their ability to model context and incorporate broader knowledge, which restricts their reasoning and generalization capabilities. Recent advances in LLMs offer strong contextual comprehension and extensive world knowledge to improve recommendation systems.
@@
Figure 1. (a) shows the composition of user interaction data. (b) and (c) illustrate previous LLM-based recommendations and our ThinkRec, respectively. (d) compares ThinkRec with baselines in three real-world datasets.
(a) shows the composition of user interaction data. (b) and (c) illustrate previous LLM-based recommendations and our ThinkRec, respectively. (d) compares ThinkRec with baselines in three real-world datasets.
Prior LLM4Rec paradigms can be separated into three categories: (1) item scoring (Liu et al., 2024; Zhang et al., 2025a), where LLMs answer binary preference questions given user and item context; (2) item generation (Li et al., 2023; Bao et al., 2024, 2025; Fu et al., 2025), which maps natural language prompts to item IDs through aligned embedding spaces or semantic identifiers; and (3) hybrid modeling (Geng et al., 2023; Zhang et al., 2024a), where a single LLM performs multiple tasks such as point-wise prediction, pair-wise ranking, or list-wise generation. Although these LLM4Rec methods differ in output formats and representation learning strategies, they fundamentally resemble System 1—the intuitive system—in cognitive science (Kahneman, 2011). They tend to match similar items based on click history, while overlooking the deeper behavioral logic. This limitation becomes evident in cases such as the one shown in Figure 1(b), where the user’s behavior over time is: dislikes “Dune”, likes “The Three-Body Problem”, and likes “Foundation” (all three are science fiction). Methods that rely on System 1-like intuition tend to infer that the user would also like “Hyperion” (science fiction) simply because it belongs to the same genre. In reality, the user might dislike philosophical or metaphysical themes in fiction, which are prominent in *Hyperion*, making it an unsuitable recommendation. Clearly, if we can leverage the vast knowledge in LLMs and fully activate their reasoning capabilities , recommendation performance can be significantly improved.
@@
This motivates our effort to push LLM4Rec from a System 1 paradigm toward a more rational, System 2-like reasoning framework. Recent concurrent works have also explored latent reasoning (Tang et al., 2025; Zhang et al., 2025b) or dual-head reasoning architectures (You et al., 2025) to address this. We raise two key questions for the advancement: 1) How to effectively balance recommendation objectives with language modeling tasks to fully exploit the reasoning capabilities of LLMs. Existing methods often prioritize direct recommendation metrics such as hit rate or ranking accuracy, overlooking LLMs’ strengths in semantic reasoning. However, blind reinforcement thinking can lead to simple next token prediction, defeating the goal of recommendation. 2) How to think more effectively in the presence of diverse user behaviors and underlying preferences. As shown in Figure 1(c), user behaviors vary widely, and uniform modeling tends to obscure personality preferences while distracting the model from salient signals, impairing its reasoning capacity. Moreover, inferring user intent solely from high-rating items and generic world knowledge limits the informational basis for accurate preference reasoning.
@@
To address these challenges, we propose Thinking-based Rec-ommendation via LLM, abbreviated as ThinkRec. One of the main problems faced is that the data and optimization goals of recommendation tasks lack the ability to activate thinking in LLMs (Challenge 1). To overcome this challenge, we designed the item augmentation and thinking activation framework for finetuning. The fine-tuned model analyzes associations in historical item information, determines user preferences, and gives explicit reasons while deriving recommendations. We extracted the metadata keywords of the items with the help of an existing summarization model as the augmentation information of the items to support the reasoning. In addition, reasoning data is synthesized using a strong reasoning model, and the reasoning capability is distilled to the local model by mixed sampling of reasoning and recommendation data. Therefore, item augmentation and thinking activation become a bridge connecting the recommendation task and the language task, making recommendations traceable. To address the difficulty of thinking diversely with rich information (Challenge 2), we add the user’s preferences (yes/no) of items to the prompts and generate personalized recommendation experts based on latent user features. In the technique, we design a dynamic Low-Rank Adaptation (LoRA) fusion method. Users are grouped by latent user features extracted with traditional recommendation models as shown in Figure 1(c). A set of base LoRAs can be fine-tuned using the grouped data and represented with corresponding user features. For each user, the engagement level of each LoRA can be determined through a gating mechanism. The difficulty of thinking is reduced through information classification and personalization mechanisms. In deployment, the reasoning ability internalized in model parameters allows ThinkRec to operate efficiently in the ranking stage: by constraining the output to a single token, the model produces batch-level results with just one forward pass.
@@
We conduct experiments on three real-world recommendation datasets, validating the rationality and effectiveness of ThinkRec. ThinkRec average outperforms state-of-the-art baselines by 7.96% in AUC and by 56.54% in METEOR. In summary, the main contributions of this work are fourfold:
  • We propose ThinkRec, a novel framework that shifts LLM-based recommendation from intuitive System 1 matching toward rational System 2-like reasoning, enabling more faithful and interpretable user preference modeling.
  • We design a data augmentation and reasoning activation framework that bridges recommendation with language modeling, leveraging metadata enrichment and reasoning guidance to trigger LLM reasoning in recommendations.
  • We introduce an instance-wise expert fusion mechanism that generates personalized recommendation experts based on latent user features, enhancing reasoning diversity and accuracy to handle diverse user behaviors.
@@
2. Related Work
2.1. LLM-based Recommendation
@@
LLM-based Recommendation. With the rapid progress of LLMs, there has been growing interest in adapting them to recommendation tasks using both textual and structured data. Early works (Bao et al., 2023; Ho et al., 2023; Lin et al., 2024a; Zhang et al., 2025a) rely on prompt-based scoring by converting recommendations into binary question-answering, while others (Ji et al., 2023a; Li et al., 2023; Lin et al., 2024b; Zhou et al., 2025) align language space with item embeddings to directly generate item IDs. More advanced approaches, such as P5 (Geng et al., 2022) and InstructRec (Zhang et al., 2024a) unify multiple sub-tasks—like rating prediction, pairwise comparison, or ranking, into a single language modeling framework. These methods demonstrate LLMs’ flexibility in expressing recommendation semantics, yet most focus on adapting output formats, treating LLMs as static scorers or selectors with limited interpretability. Beyond static prompting, recent efforts shift toward viewing LLMs as interactive agents capable of multi-step reasoning and personalized decision-making. For example, RecMind (Wang et al., 2024a) and MACRec (Wang et al., 2024b) introduce modularized reasoning structures, enabling planning-based recommendation. Others explore integrating LLMs into various stages of the pipeline—for instance, generating user or item representations (Zhang et al., 2024b) or acting as high-level decision controllers (Park and Zhang, 2025). Meanwhile, instruction tuning (Zhang et al., 2024a), chain-of-thought prompting (Liu et al., 2025) have been proposed to enhance LLMs’ semantic understanding. Despite these developments, explicit reasoning supervision and fine-grained user modeling remain underexplored. ThinkRec addresses these gaps by integrating reasoning-augmented training and personalized expert fusion, enabling user-specific recommendations grounded in structured reasoning.
@@
Reasoning Model. Recent advances have transformed LLMs from passive token predictors into structured reasoning agents by incorporating process-level supervision. Techniques such as Chain-of-Thought prompting (Wei et al., 2022), ReAct (Yao et al., 2022), and Tree-of-Thoughts (Yao et al., 2023) enable models to generate intermediate reasoning steps, improving interpretability and multistep inference quality. To further reinforce reasoning capabilities, recent works have introduced process reward models (Zhang et al., 2024c) and self-improvement pipelines using techniques like Monte Carlo Tree Search (Luo et al., 2024) and reasoning distillation (Xu et al., 2023). These trends define the emerging paradigm of Large Reasoning Models, which prioritize explicit, verifiable thinking processes over direct output. ThinkRec draws inspiration from this direction by introducing reasoning-augmented training for recommendation, aligning to embed structured reasoning into decision making.
3. Method
@@
Figure 2. Overview of the proposed ThinkRec. Keywords are extracted from the description of items with the pretrained PolyLM-Qwen-7B. Reasoning data is synthesized with a reasoning model from a sample of training data. Global LoRA is trained with total data, and base LoRAs are tuned on it with data grouping by user feature. The used LoRA, fusioned or single, is determined by the similarity between the user feature and LoRA representations.
Overview of the proposed ThinkRec.
3.1. Preliminary
@@
Integrating Collaborative Embeddings into LLMs. The sequential recommenders, such as MF (Koren et al., 2009) and LightGCN (He et al., 2020), are used for encoding collaborative information, which can be integrated into LLM embeddings. Formally, for each sample $(u,t,i_{id},l)$:
(1) $\mathbf{e}_{s}^{u}=f_{\psi}(u;\mathcal{S});\mathbf{e}_{s}^{i}=f_{\psi}(i_{id};\mathcal{S}),$
@@
(2) $\mathbf{E}_{txt}=\mathrm{WE}(\mathrm{TKZ}(i_{txt})),$
where $\mathrm{TKZ}(\cdot),\mathrm{WE}(\cdot)$ present the LLM tokenizer and word embedding layer, respectively. $\mathbf{E}_{txt}=\{\mathbf{emb}_{txt}^{1:L}\}$, where $L$ is the number of tokens and $\mathbf{emb}_{txt}^{1}\in\mathbb{R}^{1\times d_{2}}$ denotes embeddings in the language space with dimension $d_{2}$. To align collaborative embeddings $\mathbf{e}_{s}^{i},\mathbf{e}_{s}^{u}$ to LLMs’ language space, projector $proj_{\phi}(\cdot)$ is used:
(3) $\mathbf{emb}_{s}^{i}=proj_{\phi}(\mathbf{e}_{s}^{i}),\mathbf{emb}_{s}^{u}=proj_{\phi}(\mathbf{e}_{s}^{u}).$
3.2. Thinking Enhanced Recommendation
@@
3.2.1. Reasoning data construction
To support LLMs for more effective reasoning, we need more textual information in addition to item titles. Given that the raw metadata often contains excessive or noisy content, directly using long-form text may hinder the understanding and inference of LLMs. To address this, we leverage a pre-trained summarization model (Wei et al., 2023) to extract key concepts from the item metadata, and retain up to 10 representative keywords to enhance the semantic representation of each item. To model more comprehensive preference information, not just liking information, we add low-scoring interactions to the history sequence as well. Then we augment each item with title $i_{t}$, label $i_{l}\in\{yes,no\}$, feature, and keywords $i_{k}$ as follows:
$i_{txt}$ = $i_{t}$ with feature $\mathbf{emb}_{s}^{i}$ (label: $i_{l}$) with description: $i_{k}$.
Since existing recommendation datasets lack explicit reasoning traces, we synthesize a subset of samples with explainable rationale by leveraging a powerful reasoning model QwQ [^1] https://qwenlm.github.io/blog/qwq-32b/. Specifically, we sample a few thousand instances from the training data and construct textual prompts according to Appendix B.2. As in Figure 2(b), we then repeatedly query the model until a correct prediction is produced and record the latest explanation content as the reason.
3.2.2. Thinking activation
@@
(4) $\begin{array}[]{rl}\mathbf{E}^{qa}=&\mathrm{Concat}(\mathbf{E}^{q},\mathbf{E}^{a}),\\ pos=&-\mathrm{Length}(\mathbf{E}^{a}),\end{array}$
where $\mathbf{E}^{q},\mathbf{E}^{a}$ means embedding of question and answer constructing from $\mathcal{S}$, respectively. While pospos marked the position of the first answer token. This mixed sampling strategy encourages the model to jointly learn the prediction of the recommendation outcome and the semantic alignment of the reasoning behavior, within a shared representation space, as shown in Figure 2(c).
@@
(5) $\mathcal{L}_{think},logits=\mathrm{LLM}_{\theta}(\mathbf{E}^{qa}).$
@@
(6) $\begin{array}[]{rl}posid=&\mathrm{TKZ}(\mathrm{{}^{\prime}Yes^{\prime}}),\\ \hat{l}=&logits[pos][posid],\\ \mathcal{L}_{rec}=&\mathrm{BCE}(\hat{l},l).\end{array}$
To jointly optimize two objectives within a unified batch, we always use both losses simultaneously, adjusting the weight of each loss according to the instance type:
(7) $\mathcal{L}=\left\{\begin{array}[]{rl}\alpha\mathcal{L}_{rec}+\beta\mathcal{L}_{think},&thinking\\ \eta\mathcal{L}_{rec}+\gamma\mathcal{L}_{think},&recommend\end{array}\right.$.
@@
3.3. Recommendation Experts Fusion
3.3.1. Base Expert fine-tuning
To enable each expert to better capture user-specific preferences and behavioral patterns, we partition users into distinct groups based on latent features. Specifically, we utilize user embeddings derived from various pretrained small collaborative models(Lv et al., 2025a; Fu et al., 2024; Zhang et al., 2023), each of which implicitly encodes the interaction semantics of users. These embeddings serve as the basis for grouping. We aggregate the representations across all users and perform unsupervised clustering to obtain $N$ groups $\mathcal{S}^{\prime}=\{\mathcal{S}_{1:N}\}$. The resulting clusters are then used to partition the training and validating data, allowing each expert to specialize in a subset of users with similar representations to simplify the preference modeling.
We first fine-tune a global expert using all the data in the framework described in Section 3.2.2, employing LoRA-based adaptation to enable generalizable recommendations with thinking(Lv et al., 2023, 2025b; Zhan et al., 2025). Building on this global expert $\mathrm{LoRA}_{global}$, we further fine-tune the selected LoRA layers (the last 8 layers) with grouped data, enabling the model to preserve general thinking capability while adapting to finer-grained user preferences in recommendation. This produces a set of base experts $\{\mathrm{LoRA}_{1:N}\}$ that act as foundations for user-specific dynamic expert fusion.
3.3.2. Instance Wise Expert Fusion
To determine which expert is most suitable for a given user, we assign the mean of user features extracted within each group by the corresponding small model as representations of experts $\mathcal{E}=\{\mathbf{e}_{1:N}^{c}\}$. We then compute the match between user features and expert representations to estimate each base expert’s involvement in modeling the user’s preferences. The cosine similarity and softmax functions were used to obtain participation scores $\mathbf{w}^{u}=\{w^{u}_{1:N}\}$ of user $u$ based on experts:
(8) $\begin{array}[]{rl}\mathbf{e}_{n}^{c}=&\mathrm{Mean}(\mathbf{e}^{u}_{s}),u\in\mathcal{S}_{n},\\ \mathbf{w}^{u}=&\mathrm{Softmax}(\mathrm{Cosim(\mathbf{e}_{s}^{u},\mathcal{E})}/\tau),\end{array}$